An object is moving in a straight line. It initially travels at a speed of 9 meters per second, and it speeds up at a constant rate of 2 meters per second each second. Under such conditions, the distance d, in meters, that the object travels is given by the equation d = t^2 + 9t, where t is in seconds. According to this equation, how long will it take the object to travel 22 meters?
Answer by Alan3354(69443)

An object is moving in a straight line. It initially travels at a speed of 9 meters per second, and it speeds up at a constant rate of 2 meters per second each second. Under such conditions, the distance d, in meters, that the object travels is given by the equation d = t^2 + 9t, where t is in seconds. According to this equation, how long will it take the object to travel 22 meters?
--------------
d = t^2 + 9t
t^2 + 9t = 22
t^2 + 9t - 22 = 0
(t+11)*(t-2) = 0
t = 2 seconds (ignore the -11 seconds)
-------------
The 1st second it travels 10 meters.
The 2nd second it travels 12 meters.